# supplementary

### v1

@@ IV-C Memory augmentation
While experience augmentation can enhance generalization by introducing greater training diversity, the resulting variations, combined with the partially observable nature of the environment, may cause the agent to adopt over-conservative strategies as it struggles to infer the latent task context. To address this limitation, we extend augmentation to the agent’s memory by incorporating an RNN into the training framework, as illustrated in Fig. 2. The RNN, leveraging its context modeling capability, acts as an implicit task encoder. At each time step, the RNN updates its hidden state $h_{t}$ based on the observations, progressively capturing task-specific information accumulated so far. From this hidden state, the RNN generates a latent task embedding $z_{t}$, which conditions the policy, allowing it to adapt decisions based on the inferred context.
In addition to augmenting the trajectories, we extend the RNN’s hidden states to incorporate memory for the augmented tasks. As shown in Fig. 3, this is achieved by performing a forward pass with the transformed observation sequence $(o_{t}^{g})_{t=0}^{T}$ through the RNN. During this process, the RNN generates a sequence of hidden states $(h_{t}^{g})_{t=0}^{T}$ that progressively encode the evolving dynamics of the augmented observations. To ensure continuity in the context encoding, the initial hidden state $h_{0}^{g,k}$ for the current update $k$ is set to the final hidden state $h_{T}^{g,k-1}$ from the previous update. The forward pass is performed continuously throughout the episode to ensure that the RNN’s hidden states capture the cumulative dynamics of the augmented observations at each step, providing an up-to-date context encoding for subsequent updates. The latent task embeddings $(z_{t}^{g})_{t=0}^{T}$, generated from these hidden states, provide the policy with the inferred context of the augmented tasks.
This memory augmentation allows the policy to infer unobservable context information for both ID and augmented scenarios. In contrast, a simple feedforward policy without memory relies solely on immediate partial observations and lacks the ability to infer latent task context, which may lead to degraded ID performance when augmented data is introduced.
Our training process incorporating memory augmentation is outlined in Algorithm 1.
@@
V Experiments
V-A Environment and task setup
We evaluate the performance of our approach across eight legged locomotion experiments with the ANYmal D quadruped and Unitree G1 humanoid robots in Issac Lab [45]. The experiments are designed around two core task settings: position tracking and velocity tracking, with variations introduced to simulate joint failures and payload distributions:
@@ V-A Environment and task setup
We train a memory-based policy both with and without experience augmentation, referred to as Memory-Aug and Memory-ID, respectively. To assess the role of memory-based task inference, we also train a standard MLP policy with experience augmentation [30, 31], referred to as Baseline-Aug, and a standard MLP policy without augmentation, referred to as Baseline-ID.
@@
Additionally, we include randomization-based policies, where the agent is exposed to all variations mentioned above (ID + OOD) during training. These are referred to as Baseline-Rand for the MLP-only architecture and Memory-Rand for the memory-enabled architecture.
We include all implementation details in the Appendix, including observations, rewards, network architectures and hyperparameter values.
V-B Task performance and data efficiency
(a) Mean episodic returns with minimum and maximum values of policies trained with and without augmentation.
(b) Mean episodic returns with minimum and maximum values of augmentation- and randomization-based policies.
Figure 4: Evaluation of quadruped position tracking under joint failure for different policies.
Figure 5: Normalized mean episodic returns on ID and OOD tasks, with error bars indicating minimum and maximum values.
We evaluate performance using the average episodic return, calculated as the mean of undiscounted cumulative rewards across 1000 episodes. Training for each policy was conducted using five different seeds. Figure 4 presents detailed evaluations of different policies in quadruped position tracking under both ID and OOD joint failure scenarios.
As shown in Fig. 4(b), Baseline-Aug exhibits ID performance comparable to Baseline-Rand, but falls short of matching the ID performance of Baseline-ID (see Fig. 4(a)). These findings underscore how variations introduced through randomization or augmentation can cause a simple MLP policy to adopt suboptimal behaviors in partially observable settings. Lacking the ability to infer latent task context from partial observations, such policies resort to over-conservative strategies, limiting their ability to optimize performance in specific tasks.
In contrast, Memory-Aug achieves performance comparable to Memory-ID on ID tasks, as shown in Fig. 4(a). This result highlights how the memory module effectively captures task context from partial observations and utilizes this capability in experience augmentation, thereby preserving robust ID performance.
In Fig. 4(b), Memory-Aug demonstrates performance on par with Memory-Rand in both ID and OOD tasks. Notably, Memory-Aug achieves this without requiring any additional environment interactions for these OOD tasks, demonstrating significantly higher sample efficiency than Memory-Rand.
To verify that these findings are not coincidental, we evaluate the approaches across all experiments. As shown in Fig. 5, the trend of ID performance degradation in Baseline-Rand and Baseline-Aug persists, particularly in experiments involving the quadruped robot, where task diversity is more pronounced. In contrast, Memory-Rand and Memory-Aug consistently achieve comparable ID performance to Memory-ID. This again highlights the importance of the memory mechanism in partially observable settings with substantial task diversity. By leveraging its context inference capability, this memory module allows the agent to maintain robust performance in ID tasks. Furthermore, Memory-Aug achieves performance comparable to Memory-Rand in both ID and OOD tasks. However, unlike Memory-Rand, Memory-Aug eliminates the need for additional environment interactions, effectively improving sample efficiency.
V-C Behavior analysis with memory augmentation
Figure 6: Behavior analysis of different policies for quadruped position tracking under an ID KFE joint failure. The yellow marker denotes the impaired joint, the blue arrow represents the current base velocity, and the green marker indicates the goal position. Left: The Baseline-ID-trained agent actively turns sideways toward the goal within the first second, then moves laterally while dragging the impaired leg to minimize reliance on the failed joint. In contrast, agents trained with Baseline-Rand and Baseline-Aug attempt to walk directly toward the goal using very small steps. Failing to adapt to the joint failure, these agents often stumble and passively change their body orientation during tracking. Right: Agents trained with memory-enabled policies actively adjust their body orientation within the first second, effectively compensating for the impaired joint and maintaining stable locomotion.
We observe distinct strategies adopted by different policies for quadruped position tracking with impaired joint, particularly in response to HFE and KFE joint failures. Agents trained with Baseline-Rand and Baseline-Aug often attempt to walk directly toward the goal using very small steps, as shown in the left part of Fig. 6. This over-cautious behavior reflects a tendency to uniformly reduce reliance on all joints rather than identifying and compensating for the specific failed joint. Despite their careful movements, the joint impairment often leads to stumbling or passive change in body orientation during tracking.
In contrast, agents trained with Baseline-ID and Memory-ID demonstrate a more adaptive strategy under ID joint failure. They tend to first turn sideways toward the goal, then move laterally while dragging the impaired leg to minimize reliance on the failed joint (see upper row of Fig. 6). This intentional body orientation adjustment at the start enables them to maintain greater stability while effectively progressing toward the goal. Notably, agents trained with Memory-Rand and Memory-Aug exhibit similar adaptive behavior, as shown in the middle and bottom rows on the right side of Fig. 6.
To further support these findings, we conduct a quantitative analysis of feet air time, calculated as the total duration the feet remain off the ground, and feet contact frequency, measured as the total number of instances the feet make contact with the ground. These metrics provide insight into how effectively the agents adapted their motion to compensate for joint failures.
Figure 7: Motion analysis. Normalized episodic feet air time and feet contact frequency for different policies in quadruped position tracking under ID and OOD joint failures, with error bars indicating minimum and maximum values.
As illustrated in Fig. 7, in ID tasks, Baseline-Rand and Baseline-Aug demonstrate significantly shorter feet air time and more frequent ground contact compared to Baseline-ID. This aligns with our observation that agents employing these policies tend to take smaller, more cautious steps during goal tracking. This conservative behavior persists in OOD tasks, as evidenced by the low feet air time and high contact frequency of Baseline-Rand and Baseline-Aug under OOD joint failures. In contrast, Memory-Rand and Memory-Aug exhibit feet air time and contact frequency comparable to Memory-ID in ID tasks, while maintaining consistent levels in OOD tasks. This highlights the ability of Memory-Rand and Memory-Aug to effectively adapt their motion strategies in response to both ID and OOD joint failures.
We attribute this context-aware behavior to the memory-based task inference capability provided by the RNN. By leveraging past experience, the RNN enables the agent to infer the underlying task context from partial observations, including the presence and nature of joint failures. This capability enables the agent to dynamically adapt its behavior across diverse scenarios, adopting strategies that compensate for impairments and optimize task performance.
VI Hardware experiments
Figure 8: Hardware experiment. Position tracking under ID (LF HAA) and OOD (RH HAA) joint failures with Memory-Aug policy.
We tested our Memory-Aug policy in position tracking under joint failure on the ANYmal D robot. To facilitate sim-to-real transfer, we introduced randomization in terrain profiles with a noise range of $[0.0,0.05]$ during training.
As shown in Figure 8 and the supplementary video, the robot successfully tracked multiple goals consecutively, while adapting effectively to both ID and OOD joint failures. These results demonstrate the ID robustness and OOD generalization of the memory-augmented policy on real robots.
Figure 9: Hardware experiment. Position tracking under ID (LF HAA) joint failure with Baseline-Aug policy.
Additionally, we also tested Baseline-Aug policy trained under the same terrain profiles. As shown in Figure 9, the robot exhibited limited adaptability under ID joint failure, frequently losing balance during goal tracking.
VII Limitations
The main limitation of our work lies in the reliance on prior knowledge of the task structure, which is essential for augmenting the training experience to simulate diverse task conditions. This dependency restricts the selection of transformations to those that satisfy the necessary invariance assumptions, potentially limiting generalization in cases where such assumptions are difficult to establish. That said, this reliance on prior knowledge also offers the advantage of systematically utilizing our understanding of the system’s dynamics. By incorporating domain-specific insights, we can design targeted transformations that align with the inherent properties of the task. In the future, we aim to broaden the applicability of our method by exploring more flexible augmentation techniques to simulate a wider range of OOD task conditions and enhance generalization across diverse and unstructured environments.
VIII Conclusion
In this work, we have presented an RL framework that integrates task-structured experience augmentation with memory mechanisms to enhance task generalization. By simulating varied task conditions through augmented training experiences and incorporating memory-based task inference, our approach trains a single unified policy capable of handling both ID and OOD tasks without additional environment interactions.
This framework effectively addresses the limitations of policies lacking context-awareness, which tend to adopt over-conservative strategies in highly randomized and partially observable environments. Through extensive simulation experiments, we demonstrate that our approach achieves zero-shot generalization to OOD tasks while maintaining robust ID performance.
Furthermore, our approach demonstrates performance comparable to domain randomization in both ID and OOD tasks, while the latter requiring explicit environment interactions for these tasks. Notably, our method achieves this without requiring additional environment interactions for the OOD tasks, thereby overcoming the sample inefficiency associated with domain randomization.
Crucially, the memory mechanism proves essential in partially observable settings with substantial task diversity, allowing the agent to infer task context from past interactions. This facilitates zero-shot task adaptation and ensures reliable performance across both ID and OOD tasks. Moreover, our hardware experiments on a quadruped robot validate the sim-to-real transfer, demonstrating that our policy performs well on both ID and OOD tasks in real-world scenarios.
We believe this work inspires further advancements in developing adaptive RL agents capable of handling the complexities and uncertainties of real-world environments.
Appendix A Appendix
A-A Implementation details
Our experiments focus on two primary task settings:
[…]
    • ANYmal quadruped robot: The forward and lateral linear velocities are sampled uniformly between $-1.0$ and $1.0$ m/s:
    • G1 humanoid robot: The forward linear velocity is sampled uniformly between $0.0$ and $1.0$ m/s:
      $v_{x}^{\text{cmd}}\sim U([0.0,1.0]),$
      while the lateral velocity is fixed at zero:
      $v_{y}^{\text{cmd}}=0.0.$
@@ A-A Implementation details
Tables S1S6 summarize the observations, reward terms and training details for the two task settings.
@@
Remaining time $t_{\text{left}}$ 1 - -
@@
Angular roll-pitch
velocity
$\|\boldsymbol{\omega}_{xy}\|^{2}$ None -0.005
@@
Feet air time $\sum_{j=1}^{n_{\text{feet}}}(t_{\text{air},j}-t_{\text{thres}})\cdot\mathbf{1}_{\text{new-contact},j}$ 0.125 0.75
@@
Angular roll-pitch
velocity
$\|\boldsymbol{\omega}_{xy}\|^{2}$0.05 -0.05

### v2

@@ 4.3 Memory augmentation
In addition to augmenting trajectories, we extend the RNN’s hidden states to incorporate memory for the augmented tasks. As shown in Fig. 2, this is achieved by performing a forward pass with the transformed observation sequence $(o_{t}^{g})_{t=0}^{T}$ through the RNN. During this process, the RNN generates a sequence of hidden states $(h_{t}^{g})_{t=0}^{T}$ that progressively encode the evolving dynamics of the augmented observations. To ensure continuity in the context encoding, the initial hidden state $h_{0}^{g,k}$ for the current update $k$ is set to the final hidden state $h_{T}^{g,k-1}$ from the previous update. The forward pass is performed continuously throughout the episode to ensure that the RNN’s hidden states capture the cumulative dynamics of the augmented observations at each step, providing an up-to-date context encoding for subsequent updates. The resulting latent task embeddings $(z_{t}^{g})_{t=0}^{T}$ provide the policy with the inferred context of the augmented tasks. This memory augmentation allows the policy to infer unobservable context information for both ID and augmented scenarios. In contrast, a simple feedforward policy without memory cannot infer latent task context from partial observations, which may lead to degraded ID performance when augmented data is introduced.
@@
5 Experiments
5.1 Environment and task setup
We evaluate our approach across eight legged locomotion tasks using the ANYmal D quadruped and Unitree G1 humanoid robots in Isaac Lab [38]. The tasks are structured around position and velocity tracking, with variations introduced through simulated joint failures and diverse payload configurations to test both ID and OOD task generalization. We compare policies trained with (-Aug) and without (-ID) experience augmentation, using both memory-based (Memory-) and standard MLP (Baseline-) architectures. Additional baselines include randomized training over all task conditions (-Rand). Full task setup and training details are provided in the Appendix.
5.2 Task performance and data efficiency
(a) Memory-Aug avoids the ID performance degradation observed in Baseline-Aug.
(b) Memory-Aug achieves performance comparable to Memory-Rand with higher sample efficiency.
Figure 3: Evaluation of quadruped position tracking under joint failure.
Figure 4: Normalized mean episodic returns on ID and OOD tasks. Memory-Aug generalizes well to OOD tasks while preserving strong ID performance across all experiments.
Figure 3 presents ID and OOD evaluations of different policies in quadruped position tracking under joint failure scenarios. As shown in Fig. 3(b), Baseline-Aug exhibits ID performance comparable to Baseline-Rand, but falls short of matching the ID performance of Baseline-ID (see Fig. 3(a)). These findings highlight how variations introduced through randomization or augmentation can cause a simple MLP policy to adopt suboptimal behaviors in partially observable settings. Lacking the ability to infer latent task context from partial observations, such policies resort to over-conservative strategies, limiting their ability to optimize performance in specific tasks. In contrast, Memory-Aug achieves performance comparable to Memory-ID on ID tasks. This result demonstrates that the memory module effectively captures task context from partial observations and leverages it in experience augmentation, thereby preserving robust ID performance. Furthermore, Memory-Aug attains performance on par with Memory-Rand in both ID and OOD tasks. Notably, Memory-Aug achieves this without requiring any additional environment interactions for these OOD tasks, demonstrating higher sample efficiency than Memory-Rand.
Next, we evaluate our approach across different robot embodiments (quadruped and humanoid), task types (position and velocity tracking), and variations in task conditions (joint failures and payload distributions). As shown in Fig. 4, the ID performance degradation in Baseline-Rand and Baseline-Aug persists, particularly in quadruped experiments, where task diversity is more pronounced. In contrast, Memory-Rand and Memory-Aug consistently match the ID performance of Memory-ID. By inferring latent task context from past experience, the memory module enables robust ID performance, even when trained under diverse task conditions. Furthermore, Memory-Aug consistently matches the performance of Memory-Rand in both ID and OOD tasks without requiring additional environment interactions, thereby improving sample efficiency.
5.2.1 Latent task embedding analysis
Figure 5: PCA visualization of latent task embeddings $z$ (LF, in red) and $z_{g}$ (RF, LH, RH, in blue) for joint failures across different joint types (HAA, HFE and KFE). The distinct clustering of tasks suggests that the learned latent space effectively captures task-specific features.
To better understand how Memory-Aug policy captures task information, we analyze the latent task embeddings $z$ and $z_{g}$ produced by the memory module in the quadruped position tracking task under joint failures. Figure 5 presents a PCA visualization of the latent embeddings collected across various joint failure scenarios. Notably, the embeddings reveal an emergent symmetry across different legs. This spatial symmetry in the latent representation aligns closely with the physical symmetry of the robot’s morphology and the underlying task structure. This finding highlights the capacity of the memory module to infer and encode unobservable task context, allowing the policy to adapt effectively to both ID and augmented task conditions.
5.3 Behavior analysis with memory augmentation
Figure 6: Motion analysis.
We observe distinct foot movement patterns across policies in the quadruped position tracking task under joint failures. As shown in Fig. 6, Baseline-Rand and Baseline-Aug exhibit shorter air time and higher contact frequency compared to Baseline-ID under ID conditions. This indicates a cautious locomotion strategy characterized by smaller steps and frequent ground contact. The same pattern persists under OOD joint failures, indicating a lack of task-specific adaptation and a tendency toward over-conservative movements. In contrast, Memory-Rand and Memory-Aug exhibit feet air time and contact frequency comparable to Memory-ID in ID tasks while maintaining similar levels in OOD tasks. We attribute this consistency to the memory module’s ability to infer latent task context from partial observations, enabling context-aware adaptation.
In addition, we qualitatively examine how agents trained with different policies track goals. Figure 7 illustrates representative behaviors under an ID KFE joint failure. As shown on the left, agents trained with Baseline-Rand and Baseline-Aug tend to walk directly toward the goal using small, cautious steps. This conservative behavior aligns with the quantitative findings in Fig. 6, where these policies exhibited shorter foot air time and higher contact frequency. The observed strategy suggests a uniform reduction in joint usage rather than targeted compensation for the impaired joint. As a result, these agents often stumble or passively drift in orientation. In contrast, agents trained with Baseline-ID and Memory-ID adopt more adaptive behaviors. They first reorient their bodies sideways toward the goal, then move laterally while dragging the impaired leg to reduce reliance on the failed joint (see upper row of Fig. 7). This intentional body adjustment supports more stable and effective goal tracking. Notably, Memory-Rand and Memory-Aug agents demonstrate similarly adaptive behaviors, highlighting the memory module’s ability to infer latent task context, including the presence and nature of joint failures, and enable targeted adaptation.
Figure 7: Behavior analysis of different policies for quadruped position tracking under an ID KFE joint failure (yellow marker). The blue arrow indicates base velocity; the green marker shows the goal. Left: The Baseline-ID agent actively turns sideways toward the goal within the first second and moves laterally, dragging the impaired leg to reduce reliance on it. In contrast, Baseline-Rand and Baseline-Aug agents take small forward steps, often stumbling and passively drifting in orientation during tracking. Right: Memory-enabled agents actively adjust their body orientation within the first second, effectively compensating for the impaired joint and maintaining stable locomotion.
5.4 Hardware experiments
Figure 8: Hardware experiment. The Memory-Aug-trained agent successfully tracks goals under both ID (LF HAA) and OOD (RH HAA) joint failures.
We deploy our Memory-Aug policy on the ANYmal D robot for position tracking under joint failure. To facilitate sim-to-real transfer, we introduce randomization in terrain profiles with a noise range of $[0.0,0.05]$ during training. As shown in Fig. 8, the robot successfully tracks multiple goals while adapting to both ID and OOD joint failures. These results demonstrate the ID robustness and OOD generalization capability of the memory-augmented policy in real-world deployment.
6 Conclusion
In this work, we present memory augmentation, an RL framework that integrates task-structured experience augmentation with memory mechanisms to enhance task generalization. By simulating varied task conditions through augmented training experiences and incorporating memory-based task inference, our approach trains a single unified policy capable of handling both ID and OOD tasks without additional environment interactions. Our approach addresses the limitations of context-unaware policies, which tend to adopt over-conservative strategies in highly randomized and partially observable environments. Extensive simulation results show that our method achieves zero-shot generalization to OOD tasks while maintaining strong ID performance. Furthermore, our approach demonstrates performance comparable to domain randomization in both ID and OOD tasks without incurring the additional sample cost typically required to expose the agent to OOD conditions during training. Crucially, the memory mechanism proves essential in supporting reliable adaptation across diverse tasks under partial observabilities. Finally, hardware experiments on a quadruped robot demonstrate the effectiveness of our approach in real-world settings, validating sim-to-real transfer for both ID and OOD scenarios. We believe this work inspires further advancements in developing adaptive RL agents capable of handling the complexities and uncertainties of real-world environments.
Limitations
Our work relies on prior knowledge of the task structure to augment the training experience and simulate diverse task conditions. A key limitation, however, lies in the restricted set of transformations that satisfy the required invariance assumptions to ensure the safe incorporation of augmented data into the policy update. This constraint narrows the range of task variations we can simulate, limiting generalization in scenarios where suitable transformations are difficult to define or do not preserve the required properties. That said, this reliance also allows for the principled design of augmentations that are grounded in domain-specific insights and aligned with the inherent structure of the task. More flexible techniques for experience augmentation would enable simulation of a wider range of OOD task conditions and enhance generalization across diverse and unstructured environments. While this paper demonstrates the utility of experience augmentation for task generalization under simple symmetry transformations, extending its applicability to more flexible augmentation techniques is left for future work.
Appendix A Appendix
A.1 Training algorithm
Algorithm 1 outlines our training process with memory augmentation.
@@ A.1 Training algorithm
A.2 Experiment details
A.2.1 Environment and task setup
Our experiments focus on two primary task settings:
[…]
    • Quadruped robot: The forward and lateral linear velocities are sampled uniformly between $-1.0$ and $1.0$ m/s:
    • Humanoid robot: The forward linear velocity is sampled uniformly between $0.0$ and $1.0$ m/s:
      $v_{x}^{\text{cmd}}\sim U([0.0,1.0]),$
      while the lateral velocity is fixed at zero:
      $v_{y}^{\text{cmd}}=0.0.$
Additionally, we introduce variations to simulate joint failures and payload distributions:
@@ A.2.1 Environment and task setup
We train a memory-based policy both with and without experience augmentation, referred to as Memory-Aug and Memory-ID, respectively. To assess the role of memory-based task inference, we also train a standard MLP policy with experience augmentation [25, 26], referred to as Baseline-Aug, and a standard MLP policy without augmentation, referred to as Baseline-ID. Additionally, we include randomization-based policies, where the agent is exposed to all variations mentioned above (ID + OOD) during training. These are referred to as Baseline-Rand for the MLP-only architecture and Memory-Rand for the memory-enabled architecture. Training for each policy was conducted using five different seeds.
@@
Tables S1S6 summarize the observations, reward terms and training details for all task settings.
@@
Normalized remaining time $t_{\text{left}}$ 1 - -
@@
Angular roll-pitch velocity $\|\boldsymbol{\omega}_{xy}\|^{2}$ None -0.005
@@
Feet air time $\sum_{j=1}^{n_{\text{feet}}}(t_{\text{air},j}-t_{\text{thres}})\cdot\mathbf{1}_{\text{new-contact},j}$ 0.1 0.75
@@
Angular roll-pitch velocity $\|\boldsymbol{\omega}_{xy}\|^{2}$0.05 -0.05
@@ A.2.2 Motion analysis metrics
A.2.3 Additional hardware experiments
Figure S9: Position tracking under ID (LF HAA) joint failure with Baseline-Aug policy.
In addition to the Memory-Aug policy, we also test the sim-to-real transfer of the Baseline-Aug policy, trained under the same terrain profiles. As shown in Figure S9, the robot exhibits limited adaptability when subjected to an ID joint failure, often losing balance and struggling to maintain stable locomotion while tracking the target. This behavior aligns with our simulation results, where the lack of context awareness in the baseline policy limited its ability to identify and adapt to specific joint failures.